# Complex magnetic landscape in kagome metal TbTi3Bi4 and competition with family of Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals

### v1

@@ Abstract
Here we present the discovery and characterization of TbTi3Bi4, the smallest rare-earth containing member of the LnTi3Bi4 family. We previously reported on the LnTi3Bi4 (Ln: La–Gd3+, Eu2+, Yb2+) kagome metals. This structure type represents a magnetically diverse and exfoliable family of kagome metals. Detailed characterization of the magnetic and electronic transport properties of TbTi3Bi4 reveals a highly complex landscape of magnetic phases arising from an antiferromagnetic ground state. The system exhibits an intimate connection between the magnetism and magnetotransport, exhibiting sharp switching from positive (+40%) to negative magnetoresistance (-50%) near a metamagnetic transition. Unlike the Lu-Gd analogs, TbTi3Bi4 cannot be synthesized by the methodology presented previously due to phase competition with another, unreported class of materials, the Ln2-xTi6+xBi9 (x1.7–1.2) (Ln: Tb–Lu) shurikagome metals. Like the LnTi3Bi4 kagome metals, the Ln2-xTi6+xBi9 family exhibits quasi-2D networks of titanium and chains of rare-earth. We present the structures and some select basic properties of the Ln2-xTi6+xBi9 family alongside the newly discovered TbTi3Bi4. A discussion regarding the phase competition between the LnTi3Bi4 and Ln2-xTi6+xBi9 families helps frame the difficulty in synthesizing LnTi3Bi4 compounds with small Ln species and provides a strategy to circumvent formation of Ln2-xTi6+xBi9. 111Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).
@@ I Introduction
Spurred on by continued discoveries in the AM3X5, AM6X6, and CoSn-type compounds, kagome metals have proven to be a rich space for exploring the intersection of solid-state chemistry and fundamental condensed matter physics. The impetus for much of the inspiration comes from the basic electronic structure afforded by the kagome motif – one which can contain Dirac points, flat bands, and Van Hove singularities. [1, 2, 3, 4]. Derived from the triangular lattice, the kagome motif have shown remarkable structural and chemical diversity, allowing for a rich interaction between structural chemistry, electronic properties, and magnetism. As a prime example, the AM3X5 compounds have been heavily investigated for complex interactions between electronic and structural instabilities (charge density wave, superconductivity).[5, 6, 7, 8, 9, 10, 11, 12, 13] The recent discovery of CsCr3Sb5 has also included the possibility of exploring the interaction between a possible spin density wave and superconductivity.[14] In fact, embedding magnetism into the kagome motif underpins much of the research into the AM6X6 and CoSn-type compounds, which continues to provide the motivation for the development of new kagome materials.[15, 16, 17, 18, 16, 19, 20, 21, 22, 23, 24, 25, 26]
One fundamental way to develop new materials is to perturb simpler, “fundamental” prototype structures. Take, for example, the CoSn family of compounds. By adding a filler atom in the natural voids formed by the structure, the CoSn-type stacking gives way to another of the most researched kagome families, the AM6X6. Further perturbations on the AM6X6 prototype yield a grandchild of the CoSn prototype – the AM3X4 kagome metals.
@@
We reported previously on the discovery of both the antimonide AV3Sb4 (A: Eu2+, Yb2+) [27] and bismuthide ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals,[28] which make up a larger family of AM3X4 compounds.[27, 28, 29, 30, 31, 32, 33] This family boasts both magnetic and non-magnetic compounds through the diverse choice of A cations, which are typically electropositive rare-earth and alkaline-earth metals. The A-site elements comprise zig-zag chains which run parallel to the M-site kagome layers. Remarkably, unlike the CoSn and AM6X6 parent structures, the AM3X4 compounds maintain a much more quasi-2D crystal structure. Crystals are easily exfoliated (similar to the AM3X5 family), which makes them an attractive option for exploring both magnetism and the kagome band structure. Accordingly, the AM3X4 compounds have been the focus of a growing body of work, particularly studies involving angle-resolved photoemission (ARPES).[32, 34, 35, 36, 37, 38]
@@
Under the conditions presented in our previous work, LnTi3Bi4 compounds containing rare-earth atoms smaller than Gd3+ did not appear. In this work, we present a modified growth procedure which allows stabilization of TbTi3Bi4, the smallest rare-earth cation reported in the LnTi3Bi4 structure thus far. Our investigation into the smaller rare-earth species (Tb–Lu) also unveiled an intimate competition between the LnTi3Bi4 kagome metals and a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals. By understanding the phase diagram surrounding the LnTi3Bi4 compound, we are able to subvert the formation of Tb2-xTi6+xBi9 and stabilize single crystals of TbTi3Bi4. This work details our rationale for the stabilization of TbTi3Bi4 alongside a discussion of its magnetic and electronic transport properties. Remarkably, TbTi3Bi4 presents some of the most rich and complex magnetism in the known AM3X4 compounds, reinforcing the need to understand and control the formation of new phases. Our work also briefly investigates the structural and magnetic properties of the new Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals, which have strong chemical disorder and unusual magnetic behavior.
@@ II.1 Single Crystal Synthesis
TbTi3Bi4 single crystals are grown through a bismuth self-flux. Tb metal (Ames), Ti (Alfa 99.9% powder), and Bi (Alfa 99.999% low-oxide shot) were measured in a 2:3:20 ratio into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit.[39] The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules. The flux was heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h. The sample was then quenched by removing it from the furnace and immediately placing in a preheated 600°C furnace. The growth proceeded with a cooling rate of 2°C/hr to 400°C, where excess bismuth was removed through centrifugation. Ln2-xTi6+xBi9 single crystals are grown in a similar manner, except the samples are cooled linearly from 1100°C to 600°C at a rate of 2°C/hr before centrifugation at 600°C.
@@ II.2 Bulk Characterization
Single crystals were mounted on kapton loops with Paratone oil for single crystal x-ray diffraction (SCXRD). Diffraction data were collected at 100 K on a Bruker D8 Advance Quest diffractometer with a graphite monochromator using Mo K radiation ( = 0.71073 Å). Data integration, reduction, and structure solution was performed using the Bruker APEX4 software package. As noted previously, diffraction on LnTi3Bi4 crystals is challenging and requires face-indexing absorption corrections.[28] All atoms were refined with anisotropic thermal parameters. CIF files for all structures are included in the ESI. Laue diffraction utilizing a Photonic Science vertical Laue system was used to orient crystals of TbTi3Bi4 for physical property measurements.
@@
Magnetization measurements of LnTi3Bi4 single crystals were performed in a 7 T Quantum Design Magnetic Property Measurement System (MPMS3) SQUID magnetometer in vibrating-sample magnetometry (VSM) mode. Additional measurements below 1.8 K on the MPMS3 utilized an additional He3 insert. Angle-resolved magnetization measurements were performed on a 7 T Quantum Design Magnetic Property Measurement System (MPMSXL) equipped with a rotator stage.
@@
Heat capacity measurements were performed in a Quantum Design 9 T Dynacool Physical Property Measurement System (PPMS) equipped a dilution refrigerator insert. Thermal contact was achieved using Apezion N-grease. Electronic resistivity measurements were performed in a Quantum Design 9 T Dynacool PPMS. Single crystals were mounted to a sheet of sapphire and then the sapphire plate was then adhered to the sample puck stage using GE varish. The sapphire could be rotated between different orientations to achieve measurements with the magnetic field both within and out-of-plane. Sample surfaces were cleaned by removing several layers using adhesive tape, and then contacts were made using silver paint (DuPont cp4929N-100) and platinum wire (Alfa, 0.05 mm Premion 99.995%).
@@ III.1 Crystal Structure and Synthesis
Our prior work identified that the ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals are readily grown from a bismuth self-flux.[28] However, growth of ATi3Bi4 crystals containing smaller rare-earth metals (Tb–Lu) was stymied due to preferential formation of Ln2-xTi6+xBi9. Figure 1(a) provides a simple schematic showing the lanthanide row sorted in order of the 9-coordinate Shannon ionic radius. Within the ATi3Bi4 compounds, both Eu and Yb appear to preferentially adopt the 2+ oxidation state. As such, the radii of Eu2+ and Yb2+ are analogous to the larger La3+ and Pr3+, respectively. Large cations preferentially stabilize the ATi3Bi4 phase (denoted by blue), and small cations preferentially stabilize Ln2-xTi6+xBi9 (green). Within this work, we demonstrate that Tb represents a unique cross-over between the two phases, and both TbTi3Bi4 and Tb2-xTi6+xBi9 can be obtained by tuning the growth profiles.
Figure 1(b,c) details the crystal structure of both the ATi3Bi4 (b) and Ln2-xTi6+xBi9 (c) compounds. As we have discussed ATi3Bi4 extensively before, for simplicity we opt to show only half the unit cell here (the omitted half is generated through an inversion operation, space group Fmmm. ATi3Bi4 contains two key structural motifs of interest; 1) a Ti–Ti kagome network, and 2) Ln-Ln zig-zag chains. The local coordination of the Ln species is approximately 9-fold. Crystals of ATi3Bi4 are quasi-2D and can be easily exfoliated using adhesive tape. We have observed that both M-Sb and Ln-Sb terminations can be achieved.[27] The kagome network in ATi3Bi4 compounds is only slightly distorted, though the quasi-1D zig-zag chain limits the symmetry of the system to orthorhombic.
@@
The Ln2-xTi6+xBi9 compounds can be understood as a heavily alloyed form of Ti8Bi9[40], where Ln atoms preferentially occupy two of the Ti sites. If the substitution was perfect, (e.g. Ln2Ti6Bi9) the Ln atoms would create a nearly ideal 1D chain that cuts perpendicular to the Ti–Ti network. The Ln–Ln distance is closer in Ln2Ti6Bi9 (3.8Å) when compared to ATi3Bi4 (4Å)The remaining Ti atoms form a unique network made of up Ti4 tetrahedra interconnected with corner-sharing triangles. Paying homage to both the linguistic origin and coordination of the “kagome lattice,” the tessellated unit resembles a four-pointed hira-shuriken and we’ve taken to nicknaming these compounds “shurikagome” metals.
@@
The two sites that are preferentially occupied by Ln atoms possess considerably different coordination than the remaining Ti sites in the parent Ti8Bi9 structure. Considering bonds 3.2Å, Ti atoms within the shurikagome network have an approximate coordination of 10 (4 Ti–Ti bonds, 6 Ti–Bi bonds) and a polyhedral volume of approximately 50.7Å3. Sites preferentially occupied by Ln possess a coordination of 8, solely from Ln–Bi bonds, and form slightly distorted square anti-prisms. The two mixed Ln/Ti sites are also distinct from one another, and one has a substantially larger polyhedral volume (56.9Å3) than the other (49.5Å3). As a conceptually pleasing observation, the rare-earth atoms appear to exclusively occupy the 8-coordinate sites, and further preferentially occupy the largest of the mixed Ln/Ti sites. Generally speaking, the larger site is approximately 75–90% occupied by rare-earth, compared to 50–75% for the smaller site. As such, the total stoichiometry can be more precisely written as Ln2-xTi6+xBi9 (x1.7–1.2). CIF files for both TbTi3Bi4 and the Ln2-xTi6+xBi9 compounds are included in the ESI. We will discuss trends in the Ln substitution alongside the magnetic properties in a later section.

### v2

# Intricate magnetic landscape in antiferromagnetic kagome metal TbTi3Bi4 and interplay with Ln2-xTi6+xBi9 (Ln: TbLu) shurikagome metals

@@ Abstract
Here we present the discovery and characterization of the kagome metal TbTi3Bi4 in tandem with a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals. We previously reported on the growth of the LnTi3Bi4 (Ln: La–Gd3+, Eu2+, Yb2+) family, a chemically diverse and exfoliable series of kagome metals with complex and highly anisotropic magnetism. However, unlike the La–Gd analogs, TbTi3Bi4 cannot be synthesized by our previous methodology due to phase competition with Ln2-xTi6+xBi9 (x1.7–1.2). Here we discuss the phase competition between the LnTi3Bi4 and Ln2-xTi6+xBi9 families, helping to frame the difficulty in synthesizing LnTi3Bi4 compounds with small Ln species and providing a strategy to circumvent formation of Ln2-xTi6+xBi9. Detailed characterization of the magnetic and electronic transport properties on single crystals of TbTi3Bi4 reveals a highly complex landscape of magnetic phases arising from an antiferromagnetic ground state. A series of metamagnetic transitions creates at least 5 unique magnetic phase pockets, including a 1/3 and 2/3 magnetization plateau. Further, the system exhibits an intimate connection between the magnetism and magnetotransport, exhibiting sharp switching from positive (+40%) to negative magnetoresistance (–50%). Like the LnTi3Bi4 kagome metals, the Ln2-xTi6+xBi9 family exhibits quasi-2D networks of titanium and chains of rare-earth. We present the structures and some basic magnetic properties of the Ln2-xTi6+xBi9 family alongside our characterization of the newly discovered TbTi3Bi4. 111Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).
@@ I Introduction
Spurred on by continued discoveries in the AM3X5, AM6X6, and CoSn-type compounds, kagome metals have proven to be a rich space for exploring the intersection of solid-state chemistry and fundamental condensed matter physics. The impetus for studying kagome metals is derived from the basic electronic structure of a kagome toy model – which creates the potential for Dirac points, flat bands, and Van Hove singularities. [1, 2, 3, 4]. Indeed, the latent potential within the kagome motif is often unlocked by designing materials where the Fermi level (E) is aligned with these electronic features, as exemplified by the AV3Sb5 (A: K, Rb, Cs) kagome materials. In AV3Sb5 compounds, the Fermi level is intrinsically tuned to a Van Hove singularity, causing these nonmagnetic materials to be heavily investigated for complex interactions between electronic and structural instabilities (charge density wave, potential unconventional superconductivity, non-trivial topology).[5, 6, 7, 8, 9, 10, 11, 12, 13] The recent discovery of CsCr3Sb5 has also included the possibility of exploring the interaction between a possible spin density wave and superconductivity in proximity to a flat band near E.[14]
Indeed, CsCr3Sb5 exemplifies another key direction for the exploration of new kagome materials – the interaction between magnetic sublattices and the underlying kagome band structure. The development of new magnetic kagome systems (either with magnetic kagome sublattices, or magnetic supporting sublattices) has been remarkable. Consider, for example, the chemical diversity of the CoSn-family and it’s grandchildren, the AM6X6 compounds.[15, 16, 17, 18, 19, 17, 20, 21, 22, 23, 24, 25, 26, 27] The structural relationship between CoSn and the AM6X6 is an instructive one, as the two compounds are fundamentally linked. The CoSn prototype features large interstitial voids between the hexagonal plaquettes of adjacent kagome networks, which offers a unique opportunity for generating new kagome materials. By adding an electropositive filler atom in these natural voids, we can produce ordered superlattices – giving birth to the AM6X6 (HfFe6Sn6, P6/mmm) prototype.[28] The more complex ordering of the AM6X6 creates an explosion of structural diversity, as the ordered superlattice exhibits multiple stacking variations and intergrowths like ErFe6Sn6 (Amam), DyFe6Sn6 (Cmmm), and YFe6Sn6 (Immm).[29] The titular series – the AM3X4 kagome metals – also contain remnants of the AM6X6 stacking, producing a prototypical grandchild of the CoSn family.
@@
We previously reported the discovery of both the antimonide LnV3Sb4 (Ln: Eu2+, Yb2+) [30] and bismuthide LnTi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals,[31] which make up a larger family of AM3X4 compounds.[32, 31, 33, 34, 35, 36, 37] This family boasts both magnetic and non-magnetic compounds through the diverse choice of A cations, which are typically electropositive rare-earth and alkaline-earth metals. The A-site elements comprise zig-zag chains which run parallel to the M-site kagome layers. Remarkably, unlike the CoSn and AM6X6 parent structures, the AM3X4 compounds exhibit a much more quasi-2D, layered, exfoliable crystal structure. Developing quasi-2D kagome platforms is critical, as the dimensionality of the kagome network and the supporting sublattices have a profound effect on the realization of key features in the electronic structure.[38] Crystals are easily exfoliated (similar to the AM3X5 family), which makes them an attractive option for exploring both magnetism and the kagome band structure. Accordingly, the AM3X4 compounds have been the focus of a growing body of work, particularly studies involving angle-resolved photoemission spectroscopy (ARPES).[36, 39, 40, 41, 42, 43]
@@
Under the conditions presented in our previous work, LnTi3Bi4 compounds containing rare-earth atoms smaller than Gd3+ did not appear. In this work, we present a modified growth procedure which allows stabilization of TbTi3Bi4, the smallest rare-earth cation reported in the LnTi3Bi4 structure thus far. Our work suggests that TbTi3Bi4 is a relatively low temperature phase (500°C) and is suppressed through competition with a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals. By quenching fluxes from 1100°C to 600°C, we suppress the growth of Tb2-xTi6+xBi9 and encourage the resorption of any shurikagome nuclei. A subsequent cool from 600°C to 400°C encourages the growth of single crystals of TbTi3Bi4. This work details our rationale for the stabilization of TbTi3Bi4 alongside a discussion of its magnetic and electronic transport properties. Remarkably, TbTi3Bi4 presents some of the most rich and complex magnetism in the known AM3X4 compounds, reinforcing the need to understand and control the formation of new phases. Our work also briefly investigates the structural and magnetic properties of the new Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals.
@@ II.1 Single Crystal Synthesis
TbTi3Bi4 single crystals are grown through a bismuth self-flux. Tb metal (Ames), Ti powder (Alfa 99.9%), and Bi shot (Alfa 99.999% low-oxide) were placed into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit[44] at a ratio of 2:3:20 Tb:Ti:Bi. The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules. Samples were subsequently heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h. The samples were then quenched by removing them from the furnace and immediately placing them into a different furnace preheated to 600°C. The growth proceeded with a cooling rate of 1°C/hr to 400°C. Excess bismuth was removed through centrifugation at 400°C.
Ln2-xTi6+xBi9 single crystals are grown through a bismuth self-flux as well. Ln metal (Ames), Ti powder (Alfa 99.9%), and Bi shot (Alfa 99.999% low-oxide) were placed into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit[44] at a ratio of 2:3:20 Ln:Ti:Bi. The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules. Ln2-xTi6+xBi9 single crystals are stable over a wide range of fluxes (e.g., Ln:Ti:Bi of 2:3:12, 2:3:20, 1:1:3). Crystals featured here were grown at a ratio of 2:3:12, though the composition of the crystals did not change appreciably under the various flux compositions investigated here. In particular, we are currently unable to grow the idealized, fully saturated Ln2Ti6Bi9. Fluxes were heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h. The samples were then cooled linearly from 1100°C to 600°C at a rate of 2°C/hr. Excess bismuth was removed through centrifugation at 600°C.
@@ II.2 Bulk Characterization
Single crystals of both TbTi3Bi4 and Ln2-xTi6+xBi9 were mounted on kapton loops with Paratone oil for single crystal x-ray diffraction (SCXRD). Diffraction data were collected at 100 K on a Bruker D8 Advance Quest diffractometer with a graphite monochromator using Mo K radiation ( = 0.71073 Å). Data integration, reduction, and structure solution was performed using the Bruker APEX4 software package. As noted previously, diffraction on LnTi3Bi4 crystals is challenging and requires face-indexing absorption corrections.[31] All atoms were refined with anisotropic thermal parameters. CIF files for all structures are included in the ESI. A TF20 X-Ray System (LAUE-V-674-AC-GigE from Photonic Science) was used to characterize and orient single crystals for magnetotransport and magnetization measurements. Elemental analysis was carried out on as-grown crystals using a Hitachi-TM3000 microscope equipped with a Bruker Quantax 70 EDS system.
@@
Magnetization measurements (300 K – 1.8 K) on crystals of TbTi3Bi4 and Ln2-xTi6+xBi9 were performed in a 7 T Quantum Design Magnetic Property Measurement System (MPMS3) SQUID magnetometer in vibrating-sample magnetometry (VSM) mode. Additional measurements on TbTi3Bi4 (1.8 K – 0.3 K) utilized the Quantum Design iHe-3 He3 insert for the MPMS3. AC susceptibility measurements (3.5 K – 0.1 K) on Tm2-xTi6+xBi9 utilized a Quantum Design Dilution Refrigerator insert in a Quantum Design 9 T Dynacool PPMS. The lowest frequency data (100 Hz) was normalized to the MPMS3 data from 4 K to 1.8 K to create a continuous data set from 300 K to 0.1 K. Angle-resolved magnetization measurements on TbTi3Bi4 were performed on a 7 T Quantum Design Magnetic Property Measurement System (MPMSXL) equipped with a rotator stage.
A note regarding hysteresis in TbTi3Bi4; crystals show substantial magnetic hysteresis, and considerable care must be taken to account for magnetic and thermal history. While samples show negligible changes in low-field (500 Oe) measurements under field-cooled (FC) or zero-field-cooled (ZFC) conditions, field loops up to 3 T must be performed with consideration for the ZFC (virgin magnetization) curve. Additionally, subsequent measurements under varying temperatures (e.g. M vs. H curves at different temperatures) need to heat up above the transition temperature to “erase” the magnetic history. Specific details for each magnetic and transport measurement are included in the text alongside the respective figures.
@@
Electronic transport measurements on TbTi3Bi4 were performed in a Quantum Design 9 T Dynacool PPMS. Single crystals were mounted to a sheet of sapphire, which was subsequently adhered to the sample puck stage using GE varish. The sapphire could be rotated between different orientations to achieve measurements with the magnetic field both within and out-of-plane. Sample surfaces were cleaned by removing several layers using adhesive tape, and then contacts were made using silver paint (DuPont cp4929N-100) and platinum wire (Alfa, 0.05 mm Premion 99.995%) in a six-wire geometry for simultaneous measurements of the Hall effect and magnetoresistance. As alluded to in the magnetization methods, care must be taken to account for magnetic and thermal history. We are primarily focused on the magnetotransport results that are repeatable under continued field cycling, and will typically omit virgin magnetoresistance curves. In this case, we typically begin each measurement in a fully-polarized (e.g. 9 T) state. All transport measurements utilized a probe current of 8 mA.
Heat capacity measurements (300–1.8 K) on TbTi3Bi4, Tb2-xTi6+xBi9, and Er2-xTi6+xBi9 were performed in a Quantum Design 9 T Dynacool Physical Property Measurement System (PPMS). Additional heat capacity measurements (3.5–0.1 K) on TbTi3Bi4 utilized the Quantum Design Dilution Refrigerator insert for the Quantum Design 9 T Dynacool PPMS. Thermal contact was achieved using Apezion N-grease.
II.3 Electronic Structure Calculations
In an effort to describe the underlying electronic structure of the Ln2Ti6Bi9 family, we have performed first-principles all-electron density functional theory calculations within the generalized gradient approximation, as in several of our recent works.[31, 45, 46, 47, 48] To avoid 4f-electron-related complications we have chosen to study the Lu-based end-member Lu2Ti6Bi9 in a non-spin-polarized configuration, with Lu having a full 4f shell and thereby no magnetic character. We have chosen to neglect potential structural disorder by studying the fully-ordered, stoichiometric structure of Lu2Ti6Bi9. The results should be generally applicable to other lanthanides in this family, presuming that relevant measurements (such as ARPES) are performed at temperatures well above the onset of any magnetic order.[49] This is as in our previous work.[31] The linearized augmented plane-wave (LAPW) basis was employed, with an RKmax (the product of the smallest muffin-tin sphere radius and largest planewave expansion wavevector) of 8.0 and sphere radii of 2.50 Bohr for Bi and Lu were chosen, and 2.41 Bohr for Ti. Neither spin-orbit coupling nor the GGA+U approach (i.e., a ‘U’) was included.
@@ III.1 Crystal Structure and Synthesis
Our prior work identified that the LnTi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals are readily grown from a bismuth self-flux.[31] However, growth of LnTi3Bi4 crystals containing smaller rare-earth metals (Ln: Tb–Lu) was stymied due to extensive crystallization of the Ln2-xTi6+xBi9 compounds. Figure 1(a) provides a simple schematic showing the lanthanide row sorted in order of the 9-coordinate Shannon ionic radius. Within the LnTi3Bi4 compounds, both Eu and Yb appear to preferentially adopt the 2+ oxidation state. As such, the radii of Eu2+ and Yb2+ are analogous to the larger La3+ and Pr3+, respectively. Large cations preferentially stabilize the LnTi3Bi4 phase, and small cations preferentially stabilize Ln2-xTi6+xBi9. Within this work, we demonstrate that Tb represents a unique cross-over between the two phases, and both TbTi3Bi4 and Tb2-xTi6+xBi9 can be obtained by tuning the growth profiles.
Figure 1(b,c) details the crystal structure of both the LnTi3Bi4 (b) and Ln2-xTi6+xBi9 (c) compounds. Note that Figure 1(b) presents only half of the LnTi3Bi4 unit cell, with the omitted half generated by an inversion operation (space group Fmmm). LnTi3Bi4 contains two key structural motifs of interest; 1) a Ti–Ti kagome network, and 2) Ln–Ln zig-zag chains. The reduced symmetry of the unit cell is driven by the quasi-1D zig-zag chains, which necessitate a maximum of 2-fold rotational symmetry. Despite the orthorhombic structure, the kagome network is negligibly distorted (0.1Å out-of-plane buckling). Accordingly, the orthorhombic b/a lattice parameter ratio () is very nearly the nominal applicable to the hexagonal symmetry. The zig-zag chains dominate the magnetic properties of these materials, with the nearest-neighbor intra-chain Ln–Ln distance (4 Å) substantially closer than the inter-chain distances (6 Å). We typically model the rare-earth local coordination as approximately 9-fold, though the coordination is not uniform and requires a small range of bond lengths (3.3 Å– 3.6 Å) to recover the polyhedral environment shown in Figure 1(b).
@@
The Ln2-xTi6+xBi9 (P4/nmm) compounds can be understood as a heavily alloyed form of Ti8Bi9[50], where Ln atoms preferentially occupy two of the Ti sites (there are 4 unique Ti sites in Ti8Bi9). If the substitution was perfect, (e.g. Ln2Ti6Bi9) the Ln atoms would create a nearly ideal 1D chain that cuts perpendicular to the Ti–Ti network. The Ln–Ln distance is closer in Ln2Ti6Bi9 (3.8Å) when compared to LnTi3Bi4 (4Å), owing to the contraction from partial occupancy of Ti. The remaining Ti atoms form a unique network made of up Ti4 tetrahedra interconnected with corner-sharing triangles. Paying homage to both the linguistic origin and the corner-sharing coordination of the kagome lattice, the tessellated unit resembles a four-pointed hira-shuriken and we’ve taken to nicknaming these compounds “shurikagome” metals.
@@
Within Ti8Bi9, the two sites that are preferentially occupied by Ln in Ln2Ti6Bi9 (Ln4 and Ln5 in Figure 1(c)) possess considerably different coordination than the Ti that comprise the shurikagome network. Considering bonds 3.2 Å, Ti atoms within the shurikagome network have an approximate coordination of 10 (4 Ti–Ti bonds, 6 Ti–Bi bonds) and a polyhedral volume of approximately 50.7 Å3. However, the Ln4 and Ln5 sites possess a coordination of 8, entirely from Ln–Bi bonds, and form slightly distorted square anti-prisms. Ln4 and Ln5 are also distinct from each other. Ln4 sits within the plane of the shurikagome network, and has substantially larger polyhedral volume (56.9 Å3) owing to the support of the Ti–Ti scaffolding. The Ln5 site is slightly collapsed along c, leading to a correspondingly lower volume (49.5 Å3) and a marginally more distorted coordination. As a conceptually pleasing observation, the rare-earth atoms appear to exclusively occupy the 8-coordinate sites, and preferentially occupy the largest of the mixed Ln/Ti sites. From single crystal diffraction results, the larger site is approximately 75–90% occupied by rare-earth, compared to 50–75% for the smaller site. As such, the total stoichiometry can be more precisely written as Ln2-xTi6+xBi9 (x1.7–1.2). CIF files and associated crystallograpic tables for both TbTi3Bi4 and the Ln2-xTi6+xBi9 compounds are included in the ESI. We will discuss trends in the Ln substitution alongside the magnetic properties in a later section.